OpenStudy (jenessa_2413):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 9 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 15 over 20 11 over 20 6 over 20 5 over 20

satellite73 (satellite73):

what is left?

satellite73 (satellite73):

\[20-9-6=?\]

OpenStudy (jenessa_2413):

its 5

satellite73 (satellite73):

ok so \(\frac{5}{20}\)

satellite73 (satellite73):

which, on planet earth, we usually call \(\frac{1}{4}\) but no matter, \(\frac{5}{20}\) is what they want

OpenStudy (jenessa_2413):

so its 5 over 20 right?

satellite73 (satellite73):

yes